# arXiv:2603.01359v1[eess.SY] 02 Mar 2026

### v1

@@ 5 Discussion
The results obtained for the XB-2 spar (Section˜3) and the Airbus Helicopters H135 BMR blade (Section˜4) show that NExT-LF enables accurate identification of the modal parameters $\omega_{n}$, $\zeta_{n}$, and $\bm{\phi}_{n}$. In the investigated cases, NExT-LF provides more reliable identifications than NExT-ERA, both in terms of agreement with benchmark values and robustness of the extracted modal set. This aspect is particularly relevant as both approaches rely on NExT to obtain impulse response functions from output-only data. For the helicopter blade, NExT-ERA does not retrieve the full set of modes identified by NExT-LF, whereas NExT-LF returns a richer modal content within the frequency band of interest.
@@
Nevertheless, the helicopter blade case highlights a limitation shared by both NExT-based methods: The identification of $\bm{\phi}_{n}$ becomes less reliable above approximately $53~\mathrm{Hz}$. Beyond this frequency, the diagonal MAC values with respect to Mugnaini et al. (2022) drop below $0.7$, reaching $0.06$ (NExT-ERA) for Mode 10 and $0.09$ (NExT-LF) for Mode 14. This behaviour is likely driven by the distribution of excitation energy over the analysed bandwidth. The XB-2 spar is excited through broadband-limited random (white) noise using a shaker, whereas the H135 BMR blade identification is based on an AVT conducted in a laboratory environment, without externally applied broadband forcing (e.g., wind). The resulting excitation levels are markedly lower, which can prevent sufficient activation of the higher-frequency modes (here, above $\sim 50~\mathrm{Hz}$), ultimately reducing the spatial coherence of the estimated $\bm{\phi}_{n}$. The observed degradation therefore appears primarily NExT-dependent rather than LF-dependent. Consistently, the effect is more pronounced for NExT-ERA, which does not yield stable modes beyond $60~\mathrm{Hz}$ in this dataset. It should be mentioned that SSI-CVA does not exhibit the same limitation.
@@
A further element is related to instrumentation and reference selection. The XB-2 spar represents a simple cantilever configuration, for which NExT can retrieve meaningful impulse response functions using a single reference channel. For the helicopter blade, two reference channels are required, as accelerations are measured along two axes. However, the absence of torsional and axial measurements may also limit the ability of NExT-based procedures to reconstruct modes with significant torsional content. In this respect, the pitching modes are not retrieved with satisfactory correlation by NExT-LF and NExT-ERA, which is consistent with the lack of measured rotational degrees of freedom.
@@
Overall, the present results suggest that NExT-LF cab be extended from prior applications on civil structures Dessena et al. (2025) to aeronautically relevant configurations, provided that adequate attention is given to reference-channel selection and to the choice of the analysed frequency band and/or the excitation conditions adopted during testing.
@@ 6 Conclusions
In this work, the Natural Excitation Technique (NExT) is paired with the Loewner Framework (LF) to extend NExT-LF from the Operational Modal Analysis (OMA) of civil structures to that of aeronautically relevant configurations. The main outcomes of this study can be summarised as follows:
@@
  • The natural frequencies identified by NExT-LF are coherent to those obtained with the benchmarks methods, NExT with the Eigensystem Realization Algorithm (NExT-ERA) and Stochastic Subspace Identification with Canonical Variate Analsysis (SSI), and data;
  • The identification of the damping ratios exhibits a significantly larger dispersion across methods and benchmark data, with comparatively high relative errors observed for all considered modes, particularly in the helicopter blade case;
  • The mode shapes retrieved in the lower and highly excited frequency band show a high (near 1) correlation (modal assurance criterion) with the benchmark data;
@@
The first and third authors have been supported by the Madrid Government (Comunidad de Madrid, Spain) under the Multiannual Agreement with the Universidad Carlos III de Madrid (IA_aCTRl-CM-UC3M). The second author is supported by the Centro Nazionale per la Mobilità Sostenibile (MOST – Sustainable Mobility Center), Spoke 7 (Cooperative Connected and Automated Mobility and Smart Infrastructures), Work Package 4 (Resilience of Networks, Structural Health Monitoring and Asset Management).

### v2

@@ 5 Discussion
The results obtained for the XB-2 spar (Section˜3) and the Airbus Helicopters H135 BMR blade (Section˜4) show that NExT-LF enables accurate identification of modal parameters $\omega_{n}$, $\zeta_{n}$, and $\bm{\phi}_{n}$. In the investigated cases, NExT-LF provides more reliable identifications than NExT-ERA, both in terms of agreement with the benchmark values and robustness of the extracted modal set. This aspect is particularly relevant, as both approaches rely on NExT to obtain impulse response functions from output-only data. For the helicopter blade, NExT-ERA does not retrieve the full set of modes identified by NExT-LF, whereas NExT-LF returns a richer modal content within the frequency band of interest.
@@
Nevertheless, the helicopter blade case highlights a limitation shared by both NExT-based methods: The identification of $\bm{\phi}_{n}$ becomes less reliable above approximately $53~\mathrm{Hz}$. Beyond this frequency, the diagonal MAC values with respect to Mugnaini et al. (2022) drop below $0.7$, reaching $0.06$ (NExT-ERA) for Mode 10 and $0.09$ (NExT-LF) for Mode 14. This behaviour is likely driven by the distribution of excitation energy over the analysed bandwidth. The XB-2 spar is excited through broadband-limited random (white) noise using a shaker, whereas the H135 BMR blade identification is based on an AVT conducted in a laboratory environment, without externally applied broadband forcing (e.g., wind). The resulting excitation levels are markedly lower, which can prevent sufficient activation of the modes located in the upper portion of the investigated frequency band (here, above $\sim 50~\mathrm{Hz}$), ultimately reducing the spatial coherence of the estimated $\bm{\phi}_{n}$. It should be clarified that the term higher-order is used throughout this work to denote modes of higher rank within each modal family (e.g. 2nd, 3rd, etc. flapping, lagging, or pitching), not merely modes above a fixed frequency threshold. Quantitatively, the ANPSD in Figure˜9 shows that the spectral energy of near peaks areas above approximately $50~\mathrm{Hz}$ is less dominant than the neighbouring area, which is consistent with the progressive degradation of the MAC values reported in Table˜4 for modes in this frequency region. This interpretation is therefore supported by the measured spectral content, although a fully decoupled verification of excitation energy at each mode would require knowledge of the input forces, which are unavailable in OMA.
Beyond $\bm{\phi}_{n}$, the damping estimates deserve separate consideration. Across both benchmarks, $\zeta_{n}$ exhibits a markedly larger method-to-method dispersion than $\omega_{n}$. For the XB-2 spar, the first-mode damping error exceeds 5050% for both NExT-based methods, whereas for the H135 BMR blade several modes show relative deviations in excess of 3030%. Two factors contribute to this: (i) the NExT correlation estimates are variance-limited by the available record length, which biases the exponential decay envelope from which $\zeta_{n}$ is inferred, and (ii) spectral leakage in the FFT step preceding the LF fit broadens the apparent resonance peaks, coupling frequency and damping estimation errors. Possible mitigation strategies include longer acquisition windows to reduce correlation variance, zero-padding or multi-taper spectral estimation to limit leakage, and dedicated regularisation of the Loewner pencil to penalise physically implausible damping values; although the latter do not constitute a real issue since they can be easily discarded via stabilisation diagram. Recent work Zhu et al. (2025) further confirms that output-only damping identification on lightly damped aerospace structures remains an open challenge, with reported cross-method variability of a comparable order of magnitude. Therefore, importantly, the observed degradation appears to be primarily NExT-dependent rather than LF-dependent. Consistently, the effect is more pronounced for NExT-ERA, which does not yield stable modes beyond 60 Hz in this dataset. It should be mentioned that SSI-CVA does not exhibit the same limitation.
The two benchmarks investigated in this work effectively span two distinct excitation regimes, and their comparison provides direct evidence of how excitation level and signal-to-noise ratio affect the robustness of NExT-LF. For the XB-2 spar, which is excited by a controlled broadband shaker with an RMS amplitude of 0.305 g, NExT-LF retrieves all three target modes with frequency errors below 1.6% and MAC values at or near unity. For the H135 BMR blade, where the excitation is ambient and uncontrolled, the identification remains accurate for the strongly excited low-frequency modes (Modes 1–4, with the addition of 8, with frequency errors below 0.6% and MAC values above 0.97), but degrades progressively for the modes in the upper portion of the bandwidth, where both the PSD amplitudes and the ANPSD (Figure˜9) are markedly lower. In particular, the MAC values for Modes 10 and 14 fall below 0.2, and the damping dispersion across methods increases substantially. This contrast indicates that the identification quality of NExT-LF is closely tied to the effective signal-to-noise ratio at each mode: where the response is well above the noise floor, NExT-LF performs comparably to SSI-CVA; where it is not, the NExT-derived correlations become variance-limited and the subsequent LF fit inherits that uncertainty, whose effect is stronger for ERA.
@@
A further element is related to instrumentation and reference selection. The XB-2 spar represents a simple cantilever configuration, for which NExT can retrieve meaningful IRFs using a single reference channel. For the helicopter blade, two reference channels are required, as accelerations are measured along two axes. However, the absence of torsional and axial measurements may also limit the ability of NExT-based procedures to reconstruct modes with significant torsional content. In this respect, the pitching modes are not retrieved with satisfactory correlation by NExT-LF and NExT-ERA, which is consistent with the lack of measured rotational degrees of freedom. More broadly, sensor layout affects modal observability in two respects. First, if the measurement points are sparse or lie close to the nodal lines of a given mode, the corresponding modal amplitude at those locations is small, degrading the signal-to-noise ratio and, consequently, the correlation estimate from which the mode shape is derived. This effect is particularly relevant for the higher-order flapping and lagging modes of the H135 BMR blade. Second, closely spaced or coupled modes can share similar spectral signatures, so that the identification algorithm must resolve small frequency separations under limited spectral resolution, which increases uncertainty in both damping and mode shape estimates.
As a limitation of this study, it is worth considering that the AVT data used here is not fully representative of the BMR blade operating conditions, which are in rotating conditions while attached to the rotor hub. Nevertheless, this is a common approach in the literature; see Agneni et al. (2010) for a similar test on an Agusta-Bell AB 204 blade. Furthermore, the individual accelerometers used in the test have masses in the range of 1-7 g. Given that the total mass of a single blade assembly is approximately 40 kg Bansemir and Emmerling (1999), the combined sensor mass (below 60 g for all nine accelerometers) is less than 0.15% of the blade mass, which rules out significant mass-loading effects on the identified modal parameters. Finally, this aspect is not critical for the main scope of this work, as the H135 BMR blade dataset is employed as a benchmark for comparing NExT-LF against NExT-ERA and SSI-CVA under identical instrumentation conditions, rather than as an absolute characterisation of the blade dynamic properties.
Several practical strategies may improve the identification of higher-order modes in future campaigns. From an instrumentation standpoint, a denser sensor grid, particularly in the outer blade span, where higher-order characteristic displacements might be shorter, would increase modal observability. This would reduce the likelihood of measurement points coinciding with nodal regions. Multi-setup measurement schemes, in which overlapping sensor configurations are combined, offer an alternative when the available channel count is limited Brincker and Ventura (2015), e.g. roving. On the excitation side, broader-band or higher-amplitude forcing (e.g. impact hammer sweeps or shaker excitation) would raise the signal-to-noise ratio in the upper frequency range, while longer acquisition records would reduce the variance of the NExT correlation estimates and sharpen the spectral peaks used in the LF fit. However, this would defeat the point of an OMA campaign.
@@
Considering these findings, the results of this article demonstrate that NExT-LF can be extended from prior applications on civil structures Dessena et al. (2025) to aeronautically relevant configurations, provided that adequate attention is given to reference-channel selection and to the choice of the analysed frequency band and/or the excitation conditions adopted during testing.
@@ 6 Conclusions
In this work, the Natural Excitation Technique (NExT) is paired with the Loewner Framework (LF) to extend NExT-LF from Operational Modal Analysis (OMA) of civil structures to that of aeronautically relevant configurations.
In conclusion, the main outcomes of this study can be summarised as follows:
@@
  • The natural frequencies identified by NExT-LF are coherent to those obtained with two benchmark methods, NExT with the Eigensystem Realization Algorithm (NExT-ERA) and Stochastic Subspace Identification with Canonical Variate Analysis (SSI), and data;
  • Damping ratio estimates exhibit a significantly larger inter-method dispersion than natural frequencies and mode shapes, attributed primarily to the sensitivity of the NExT-derived correlation estimates to record length and noise level;
  • The mode shapes retrieved in the lower and highly excited frequency band show a high (near 1) correlation (modal assurance criterion) with the reference data;
@@
Future work will focus on (i) extending the applicability of NExT-LF to poorly excited higher-frequency modes, (ii) improving damping estimation through longer acquisitions and advanced spectral preprocessing, (iii) systematically characterising the robustness of NExT-LF to measurement noise via controlled noise-injection studies, and (iv) applying NExT-LF under more representative operating conditions for helicopter blades, including in-flight or ground-run scenarios. From a practical standpoint, the computational efficiency of the LF and its compatibility with output-only data make NExT-LF a candidate tool for the modal characterisation of lightweight aerospace structures in settings where controlled excitation is impractical, such as ground vibration testing of large assemblies or in-service monitoring. In such contexts, NExT-LF could support condition-aware maintenance programmes and provide input data for finite element model updating.
@@
The first author is supported by grant JDC2024-055593-I funded by MICIU/AEI/10.13039/501100011033 and by ESF+. This work has been supported by the Madrid Government (Comunidad de Madrid, Spain) under the Multiannual Agreement with UC3M (IA_aCTRl-CM-UC3M). The second author is supported by the Centro Nazionale per la Mobilità Sostenibile (MOST – Sustainable Mobility Center), Spoke 7 (Cooperative Connected and Automated Mobility and Smart Infrastructures), Work Package 4 (Resilience of Networks, Structural Health Monitoring and Asset Management).